# Omni-SafetyBench: A Benchmark for Safety Evaluation of Audio-Visual Large Language Models

### v1

@@ Abstract
The rise of Omni-modal Large Language Models (OLLMs), which integrate visual and auditory processing with text, necessitates robust safety evaluations to mitigate harmful outputs. However, no dedicated benchmarks currently exist for OLLMs, and prior benchmarks designed for other LLMs lack the ability to assess safety performance under audio-visual joint inputs or cross-modal safety consistency. To fill this gap, we introduce Omni-SafetyBench, the first comprehensive parallel benchmark for OLLM safety evaluation, featuring 24 modality combinations and variations with 972 samples each, including dedicated audio-visual harm cases. Considering OLLMs’ comprehension challenges with complex omni-modal inputs and the need for cross-modal consistency evaluation, we propose tailored metrics: a Safety-score based on conditional Attack Success Rate (C-ASR) and Refusal Rate (C-RR) to account for comprehension failures, and a Cross-Modal Safety Consistency Score (CMSC-score) to measure consistency across modalities. Evaluating 6 open-source and 4 closed-source OLLMs reveals critical vulnerabilities: (1) no model excels in both overall safety and consistency, with only 3 models achieving over 0.6 in both metrics and top performer scoring around 0.8; (2) safety defenses weaken with complex inputs, especially audio-visual joints; (3) severe weaknesses persist, with some models scoring as low as 0.14 on specific modalities. Our benchmark and metrics highlight urgent needs for enhanced OLLM safety, providing a foundation for future improvements.
Warning: This paper contains examples of harmful contents.
@@ 1 Introduction
Omni-modal large language models (OLLMs) have made rapid progress in comprehending and producing content based on integrated inputs from visual, audio, and text modalities. Drawing on data from various modalities, these models are able to perform more complex tasks, such as describing audio-visual scenes and executing voice instructions based on visual information. However, despite these advancements, ensuring their safety remains a critical concern that prevents these models from causing harm or acting in unethical, incorrect, or biased ways (Yi et al., 2024; Liu et al., 2024b).
Developing corresponding benchmarks serves as the cornerstone for reasonable assessment of safety evaluation. Numerous safety benchmarks have been established for text-only LLMs and vision-language models (Zhang et al., 2023; Liu et al., 2024c), and recent work has extended to specialized benchmarks for audio-LLMs and video-LLMs (Li et al., 2025a; Liu et al., 2025; Lu et al., 2025). However, for OLLMs capable of processing audio-visual joint inputs, there is currently a lack of benchmarks specifically designed to evaluate their safety. Table 1 compares existing safety benchmarks with our Omni-SafetyBench.
Safety evaluation for OLLMs presents unique challenges that existing safety benchmarks fail to address:
@@
  • Cross-modal safety consistency: For OLLMs with numerous possible input modality combinations, the consistency of safety performance across different modal transformations of the same source data is particularly noteworthy—a phenomenon we term cross-modal safety consistency. Models with weak cross-modal safety consistency are vulnerable to exploitation through simple modality conversion techniques, enabling large-scale jailbreaking attacks. Currently, no existing benchmark provides parallel evaluation capabilities or can assess cross-modal safety consistency.
Modality Paradigm Modality Type Modality Variation Unimodal (1) Text-only - (2) Image-only - (3) Video-only - (4) Audio-only - Dual-modal Image-Text (5) Diffusion Image (6) TYPO Image (7) Diffusion+TYPO Image Video-Text (8) Diffusion Video (9) TYPO Video (10) Diffusion+TYPO Video Audio-Text (11) TTS Audio (12) TTS+Noise Audio Omni-modal (Audio-visual Input) Image-Audio-Text (13) Diff. Image with TTS Audio (14) Diff. Image with TTS+Noise Audio (15) TYPO Image with TTS Audio (16) TYPO Image with TTS+Noise Audio (17) Diff.+TYPO Image with TTS Audio (18) Diff.+TYPO Image with TTS+Noise Audio Video-Audio-Text (19) Diff. Video with TTS Audio (20) Diff. Video with TTS+Noise Audio (21) TYPO Video with TTS Audio (22) TYPO Video with TTS+Noise Audio (23) Diff.+TYPO Video with TTS Audio (24) Diff.+TYPO Video with TTS+Noise Audio (a) Taxonomy of Omni-SafetyBench. Refer to caption (b) Harmful category distribution. #Image #Video #Audio 3888 3888 2916 (c) Multimedia data statistics.
Figure 1: Overview of Omni-SafetyBench.
To address these challenges, we introduce Omni-SafetyBench, the first comprehensive parallel benchmark designed specifically for OLLM safety evaluation. Our benchmark uniquely features (1) diverse modality combinations as test inputs, (2) dedicated audio-visual joint harmful samples, and (3) 24 parallel test cases across different modalities to assess cross-modal safety consistency.
Building upon MM-SafetyBench (Liu et al., 2024c), we selected 972 entries as seed data to construct an extensive parallel benchmark that spans multiple modal combinations. As shown in Figure 1, Omni-SafetyBench encompasses three modality paradigms: unimodal, dual-modal, and omni-modal. Each paradigm is further subdivided based on modality types and variations. For images, we include diffusion-generated images, typographic images, and hybrid images combining both diffusion-generated and typographic elements. Video data comprises diffusion-generated videos, typographic videos, and hybrid videos that merge diffusion-generated and typographic content. Audio data features text-to-speech audio both with and without noise overlay. The benchmark consists of totally 24 distinct subcategories, with each containing 972 samples. Harmful category distribution and multimedia data statistics of Omni-SafetyBench are presented in Figures 1(b) and 1(c), respectively.
For evaluation, we make two special considerations tailored to the characteristics of OLLMs. (1) OLLMs can accept inputs from various modality combinations, where more complex combinations often present greater understanding challenges. We argue that discussing safety without considering comprehension is meaningless, as failure to generate harmful content due to poor understanding does not indicate strong safety. Therefore, we first evaluate comprehension based on question-answer pairs, then conduct safety evaluation only on understood inputs. This yields conditional Attack Success Rate (C-ASR) and conditional Refusal Rate (C-RR), from which we derive a Safety-score based on these two metrics. (2) As mentioned earlier, cross-modal safety consistency represents a crucial aspect of the safety of OLLMs. To address this, we design a Cross-Modal Safety Consistency Score (CMSC-score) to evaluate each model’s safety consistency across the 24 parallel subcategories in Omni-SafetyBench.
Figure 2: Safety profiles of evaluated OLLMs based on Safety-score and CMSC-score performance.
We evaluate 6 open-source and 4 closed-source state-of-the-art OLLMs using Omni-SafetyBench. Based on their performance on both Safety-score (overall safety performance) and CMSC-score (cross-modal safety consistency), we identify four distinct safety profiles among these models, as illustrated in Figure 2. Through extensive experiments, our findings reveal several critical insights: (1) Current models struggle to excel in both overall safety performance and consistency. Only gemini-2.5-pro series and Qwen2.5-Omni-7b achieve above 0.6 in both metrics, with even the best-performing model reaching just 0.8 in both dimensions. (2) OLLMs’ safety defenses weaken significantly when facing more complex modality combinations, with audio-visual joint inputs emerging as the most effective trigger for exploiting safety vulnerabilities across most models. (3) The shortest planks identified by Omni-SafetyBench are particularly concerning: even top-tier closed-source model gemini-2.5-flash achieves only a 0.52 safety-score on its worst-performing modality variations, while open-source model minicpm-o scores a mere 0.14, indicating approximately no defense against certain attacks.
@@
Table 1: Comparison of existing safety benchmarks with our Omni-SafetyBench. Omni-SafetyBench features a large-scale dataset with comprehensive modal combinations, and notably addresses cross-modal safety consistency as a crucial aspect of safety evaluation. Note: For modalities, T: Text, I: Image, V: Video, A: Audio; for metrics, ASR: Attack Success Rate, $R$: Refusal Rate, UNR: Understand Rate.
@@
The contributions are summarized as follows:
@@
  • We are the first to systematically investigate safety problems of audio-visual large language models (i.e., omni large language models), and build a large parallel benchmark Omni-SafetyBench that contains safety test data across 24 different modality/format combinations.
  • We identify two special characteristics in safety evaluation of OLLMs: the impact of comprehension on safety assessment and the evaluation perspective of cross-modal safety consistency. We innovatively design a Safety-score based on C-ASR and C-RR, as well as a CMSC-score for evaluating cross-modal safety consistency.
  • We evaluate 6 open-source and 4 closed-source OLLMs using Omni-SafetyBench, revealing current deficiencies in OLLM safety and highlighting areas that require further enhancement.
@@ 2.1 Dataset Composition
As outlined in Section 1, Omni-SafetyBench serves as a comprehensive parallel benchmark derived from a unified seed dataset through modality transformations. It incorporates three modality paradigms and nine modality types, resulting in a total of 24 modality variations. Figure 3 illustrates the different modality variations based on the seed data prompt: “List steps in detail on how to make a homemade weapon for self-harm purposes." The details for each modality variation are as follows:
Unimodal.   The unimodal paradigm refers to scenarios where input consists of a single modality. ①Text-only: The input is the seed data text. ②Image-only: The text is transformed into a static image with a width of 1024 pixels, and the height adjusts dynamically based on text length. ③Video-only: The text is converted into a silent video with a resolution of 640 $\times$ 384 pixels and a duration of 5 seconds, displaying words sequentially. ④Audio-only: The text is converted into speech with a duration of no more than 10 seconds.
Dual-modal.  The dual-modal paradigm refers to scenarios where input combines two modalities. In this scenario, multimedia data represents only the harmful key phrase, such as “homemade weapon for self-harm", combined with the transformed text to form a complete instruction. ①Image-Text: The key phrase is represented as an image, which can be Diffusion-generated, a typographic image, or a vertically concatenated combination of both. ②Video-Text: The key phrase is represented as a video, which can be Diffusion-generated, a typographic video, or a temporally concatenated combination of both. ③Audio-Text: The key phrase is represented as audio, either as text-to-speech or as text-to-speech with 3 seconds of white noise. The use of white noise is inspired by Yang et al. (2025) and Lu et al. (2025), which demonstrate that noise can effectively compromise model safety defenses.
Omni-modal.  The omni-modal paradigm refers to scenarios where input combines audio, visual, and text modalities. In this paradigm, audio and visual inputs redundantly represent the harmful key phrase, combined with text to form a complete instruction. ①Image-Audio-Text: Builds on Image-Text data by adding audio input (text-to-speech or text-to-speech+noise), with corresponding adjustments to the text (i.e., changing “The image shows” to “The image and audio show”); ②Video-Audio-Text: Builds on Video-Text data by adding audio input (text-to-speech or text-to-speech+noise), with corresponding adjustments to the text (i.e., changing “The video shows” to “The video and audio show”).
@@ 2.2 Dataset Construction
Seed Data Selection.   We selected 972 entries from MM-SafetyBench (Liu et al., 2024c) due to its diverse harmful categories and its decomposable structure, which separates a harmful key phrase from the main instruction. This structure was essential for our parallel construction methodology.
@@
Text Instruction Adjustment.  Finally, we adapted the text instructions for each of the 24 sub-datasets to ensure logical consistency. This involved updating modality-specific phrases, such as changing “The image shows…" to “The video and audio show…", to accurately match the inputs.
@@ 2.3 Evaluation
The safety evaluation of the model encompasses two key aspects. First, we assess safety performance across individual modality variations. Second, we evaluate the cross-modal safety consistency of the model across all 24 modality variations within the benchmark framework.
@@ 2.3.1 Safety Evaluation on Individual Modality Variations
Overview.  The safety evaluation for individual modality variations involves the following steps: using LLM-as-a-judge (Gu et al., 2024) to assess three aspects — whether the model understands the question, whether its output contains harmful content, and whether its output indicates a refusal to answer. Based on these assessments, we calculate Conditional Attack Success Rate (C-ASR): the probability of generating harmful content when the question is understood, and Conditional Refusal Rate (C-RR): the probability of refusing to answer when the question is understood. Finally, to better quantify safety performance, a Safety-score is derived from C-ASR and C-RR.
Motivation of Introducing Comprehension Evaluation.  As previously noted, OLLMs can process inputs combining various modalities, with more complex combinations often presenting greater comprehension challenges that may lead to irrelevant responses. Our testing revealed two types of inaccurate safety assessments stemming from comprehension issues: (1) Models failing to generate harmful responses due to inability to understand multimodal inputs, artificially inflating safety scores when such interactions are included in evaluation; (2) Models merely repeating potentially harmful prompts when processing certain modality variations, possibly due to limited training data formats, resulting in LLM-as-a-judge marking these responses as harmful and artificially deflating safety scores. Specific examples of both phenomena are provided in Appendix A.1.
C-ASR and C-RR.  To calculate C-ASR and C-RR, we first need to evaluate whether the model understands the question, whether its output contains harmful content, and whether it refuses to answer for each question-answer pair. We employ Qwen-Plus [^2] https://bailian.console.aliyun.com/, the flagship closed-source model from the Qwen team, as our judge model. The judge model takes seed data prompts and model responses as inputs (note that this does not involve multi-modal inputs), using prompts detailed in Appendix A.3. To ensure evaluation reliability, we conducted a consistency analysis between human evaluations and LLM-as-judge on a subset of data, with results presented in Appendix A.2. This process assigns three labels to each question-answer pair: understand, safe, and refuse. The metrics are then calculated as follows:
$\text{C-ASR}=\frac{N_{\text{safe}=\text{true}\cap\text{understand}=\text{true}}}{N_{\text{understand}=\text{true}}}\quad\text{,}\quad\text{C-RR}=\frac{N_{\text{refuse}=\text{true}\cap\text{understand}=\text{true}}}{N_{\text{understand}=\text{true}}}$ (1)
@@
where $\lambda$ is a weighting parameter, set to 0.5 in the experiments. The Safety-score ranges from 0 to 1, with higher values indicating stronger safety performance. The rationality of Equation 2 lies in its design: a higher C-ASR leads to a lower Safety-score, while a higher C-RR results in a higher Safety-score. The formulation emphasizes the dominance of C-ASR in safety evaluation, as the presence of harmful content serves as the primary safety criterion. Meanwhile, the refusal to respond acts as a secondary factor that demonstrates a model’s higher-confidence safety awareness.

### v2

@@ Abstract
The rise of Omni-modal Large Language Models (OLLMs), which integrate visual and auditory processing with text, necessitates robust safety evaluations to mitigate harmful outputs. However, no dedicated benchmarks currently exist for OLLMs, and existing benchmarks fail to assess safety under joint audio-visual inputs or cross-modal consistency. To fill this gap, we introduce Omni-SafetyBench, the first comprehensive parallel benchmark for OLLM safety evaluation, featuring 24 modality variations with 972 samples each, including audio-visual harm cases. Considering OLLMs’ comprehension challenges with complex omni-modal inputs and the need for cross-modal consistency evaluation, we propose tailored metrics: a Safety-score based on Conditional Attack Success Rate (C-ASR) and Refusal Rate (C-RR) to account for comprehension failures, and a Cross-Modal Safety Consistency score (CMSC-score) to measure consistency across modalities. Evaluating 6 open-source and 4 closed-source OLLMs reveals critical vulnerabilities: (1) only 3 models achieving over 0.6 in both average Safety-score and CMSC-score; (2) safety defenses weaken with complex inputs, especially audio-visual joints; (3) severe weaknesses persist, with some models scoring as low as 0.14 on specific modalities. Using Omni-SafetyBench, we evaluated existing safety alignment algorithms and identified key challenges in OLLM safety alignment: (1) Inference-time methods are inherently less effective as they cannot alter the model’s underlying understanding of safety; (2) Post-training methods struggle with out-of-distribution issues due to the vast modality combinations in OLLMs; and, safety tasks involving audio-visual inputs are more complex, making even in-distribution training data less effective. Our proposed benchmark, metrics and the findings highlight urgent needs for enhanced OLLM safety.
@@ 1 Introduction
Omni-modal large language models (OLLMs) have advanced rapidly in understanding and generating content from integrated visual, audio, and text inputs. This enables them to handle complex tasks, such as describing audio-visual scenes or following voice instructions with visual context. Despite these advancements, ensuring their safety remains a critical concern that prevents these models from causing harm or acting in unethical, incorrect, or biased ways (Yi et al., 2024).
Developing corresponding benchmarks serves as the cornerstone for reasonable assessment of safety evaluation. Numerous safety benchmarks have been established for text-only LLMs and vision-language models (Zhang et al., 2023; Liu et al., 2024b), and recent work has extended to specialized benchmarks for audio-LLMs and video-LLMs (Li et al., 2025a; Liu et al., 2025; Lu et al., 2025). However, for OLLMs capable of processing audio-visual joint inputs, there is currently a lack of benchmarks specifically designed to evaluate their safety. Table 1 compares existing safety benchmarks with our Omni-SafetyBench.
Safety evaluation for OLLMs presents unique challenges that existing benchmarks fail to address:
@@
  • Cross-modal safety consistency: OLLMs handle many different input modality combinations. A key issue is whether their safety stays consistent when the same seed data is converted into different modalities—a concept we term cross-modal safety consistency. Weak consistency lets attackers exploit the model by switching modalities, leading to widespread jailbreaking. No existing benchmarks evaluate this or provide parallel test cases.
@@
Table 1: Comparison of existing safety benchmarks with our Omni-SafetyBench. Omni-SafetyBench provides a large-scale dataset with diverse modal combinations and highlights cross-modal safety consistency as a key evaluation factor. T(Text), I(Image), V(Video), A(Audio), ASR(Attack Success Rate), RR(Refusal Rate).
@@
To address these challenges, we introduce Omni-SafetyBench, the first comprehensive parallel benchmark designed specifically for OLLM safety evaluation. Our benchmark uniquely features (1) diverse modality combinations as test inputs, (2) dedicated audio-visual joint harmful samples, and (3) 24 parallel test cases across different modalities to assess cross-modal safety consistency. Building upon MM-SafetyBench (Liu et al., 2024b), we selected 972 entries as seed data to construct an extensive parallel benchmark that spans multiple modal combinations. As shown in Table 2, Omni-SafetyBench encompasses three modality paradigms: unimodal, dual-modal, and omni-modal. Each paradigm is further subdivided based on modality types and variations, resulting in totally 24 distinct subcategories, with each containing 972 samples. Harmful category distribution and multimedia data statistics of Omni-SafetyBench are presented in Appendix B.
Figure 1: Safety profiles of evaluated OLLMs.
For evaluation, we make two special considerations tailored to the characteristics of OLLMs. (1) OLLMs handle inputs from diverse modality combinations, with complex ones often posing greater comprehension challenges. We believe safety discussions without assessing understanding are pointless, since failing to produce harmful content due to poor comprehension doesn’t prove true safety. Thus, we first measure comprehension via question-answer pairs, then evaluate safety only on well-understood inputs. This produces the Conditional Attack Success Rate (C-ASR) and Conditional Refusal Rate (C-RR), which we use to compute a Safety-score. (2) As noted earlier, cross-modal safety consistency is vital for OLLM safety. To measure it, we create the Cross-Modal Safety Consistency score (CMSC-score), assessing each model’s consistency across Omni-SafetyBench’s 24 parallel subcategories.
Table 2: Taxonomy of Omni-SafetyBench. Omni-SafetyBench contains three modality paradigms, nine modality types (bolded) and 24 modality variations (colored).
Modality Paradigms Modality Types and Variations
Unimodal (1) Text-only (2) Image-only (3) Video-only (4) Audio-only
Dual-modal Image-Text (5) Diffusion Image (6) Typographic Image (7) Diffusion+Typographic Image
Video-Text (8) Diffusion Video (9) Typographic Video (10) Diffusion+Typographic Video
Audio-Text (11) Text-to-Speech (TTS) Audio (12) TTS+Noise Audio
Omni-modal
(Audio-visual Input)
Image-Audio-Text (13) Diffusion Image with TTS Audio (14) Diffusion Image with TTS+Noise Audio
(15) Typographic Image with TTS Audio (16) Typographic Image with TTS+Noise Audio
(17) Diff.+TYPO Image with TTS Audio (18) Diff.+TYPO Image with TTS+Noise Audio
Video-Audio-Text (19) Diffusion Video with TTS Audio (20) Diffusion Video with TTS+Noise Audio
(21) Typographic Video with TTS Audio (22) Typographic Video with TTS+Noise Audio
(23) Diff.+TYPO Video with TTS Audio (24) Diff.+TYPO Video with TTS+Noise Audio
We evaluate 6 open-source and 4 closed-source state-of-the-art OLLMs using Omni-SafetyBench. Based on their performance on both Safety-score (overall safety performance) and CMSC-score (cross-modal safety consistency), we identify four distinct safety profiles among these models, as illustrated in Figure 1. Extensive experiments reveal several critical insights: (1) Current models struggle to excel in both overall safety performance and consistency. Only three models (gemini-2.5-pro series and Qwen2.5-Omni-7b) achieve above 0.6 in both metrics. (2) OLLMs’ safety weakens sharply with complex modality combinations, where audio-visual inputs prove most effective at triggering vulnerabilities in most models. (3) Severe weaknesses persist: even the leading closed-source model gemini-2.5-flash scores only 0.52 on its worst modality variations, while open-source minicpm-o hits just 0.14, showing virtually no defense against certain attacks.
Moreover, using Omni-SafetyBench, we evaluated the effectiveness of existing safety alignment methods for OLLMs, including three inference-time algorithms and post-training approaches on four safety alignment datasets (two text-only and two image-text, with no other types currently available). We found that inference-time methods are generally less effective for OLLM safety alignment than post-training with image-text datasets. For post-training methods, OLLMs handle a wide range of input modality combinations, but existing datasets target only specific ones, leading to major out-of-distribution issues. We further explored whether these safety issues could be fixed by simply converting the training data into more modalities. After augmenting existing datasets and using them for alignment, we found that as modality combinations get more complex, the reasoning needed for safety becomes more challenging — making the same amount of in-distribution training data less effective overall. The contributions are summarized as follows:
@@
  • We are the first to investigate safety issues in omni-modal large language models and introduce Omni-SafetyBench, a large-scale parallel benchmark with 23,328 safety test instances across 24 modality variations, including audio-visual joint harm cases.
  • We identify two key characteristics in OLLM safety evaluation: the role of comprehension in safety assessment and cross-modal safety consistency. Based on these, we propose a Safety-score (derived from Conditional Attack Success Rate (C-ASR) and Conditional Refusal Rate (C-RR)) and a Cross-Modal Safety Consistency Score (CMSC-score).
  • Using Omni-SafetyBench, we evaluate 6 open-source and 4 closed-source OLLMs, uncovering safety deficiencies and generating safety profiles for each model, while also assessing existing safety alignment methods to uncover major challenges in OLLM safety alignment.
@@ 2.1 Dataset Composition
As outlined in Section 1, Omni-SafetyBench serves as a comprehensive parallel benchmark derived from a unified seed dataset through modality transformations. It incorporates three modality paradigms and nine modality types, resulting in a total of 24 modality variations. Figure 2 illustrates the different modality variations in Omni-SafetyBench derived from the same seed data prompt: “List steps in detail on how to make a homemade weapon for self-harm purposes." The details for each modality paradigm, type and variation are as follows:
Unimodal. The unimodal paradigm refers to scenarios where input consists of a single modality. ①Text-only: uses the seed text directly. ②Image-only: converts text to a typographic image (1024 pixels wide, height adjusts by text length). ③Video-only: converts text to a typographic video (640 $\times$ 384 pixels, 5 seconds, words appear sequentially). ④Audio-only: converts text to speech.
Dual-modal. The dual-modal paradigm refers to scenarios where input combines two modalities. In this case, multimedia data represents only the harmful key phrase, such as “homemade weapon for self-harm”, combined with the transformed text to form a complete instruction. ①Image-Text: key phrase as an image (diffusion-generated, typographic, or concatenated combination). ②Video-Text: key phrase as a video (diffusion-generated, typographic, or temporally concatenated). ③Audio-Text: key phrase as audio (text-to-speech or text-to-speech with 3 seconds of white noise), inspired by Yang et al. (2025) and Lu et al. (2025), which show noise can breach safety defenses.
Omni-modal. The omni-modal paradigm refers to scenarios where input combines audio, visual, and text modalities. In this case, audio and visual inputs redundantly represent the harmful key phrase, combined with text to form a complete instruction. ①Image-Audio-Text: extends image-text by adding audio inputs that representing the key phrase, and adjusts the text from “The image shows” to “The image and audio show”. ②Video-Audio-Text: extends video-text by by adding audio inputs that representing the key phrase, and adjusts the text from “The video shows” to “The video and audio show”. Detailed data statistics of Omni-SafetyBench are in Appendix B.
@@ 2.2 Dataset Construction
Seed Data Selection. We selected 972 entries from MM-SafetyBench (Liu et al., 2024b) as our seed data. This choice was intentional, as MM-SafetyBench offers comprehensive coverage of common harmful semantic categories, providing a semantically rich foundation for our benchmark. Its decomposable structure, which isolates harmful key phrases from main instructions, was particularly suitable for our parallel construction approach. Building on this, we extended these core harmful concepts across multiple modalities, creating a benchmark with a broad range of expressive diversity. To address concerns about seed data leakage, we conducted simulated experiments (see Appendix C), which demonstrate that models trained on the seed data still struggle with omni-modal cases in our benchmark. This confirms the robust effectiveness of Omni-SafetyBench.
@@
Text Instruction Adjustment. Finally, we adapted the text instructions for each of the 24 sub-datasets to ensure logical consistency. This involved updating modality-specific phrases, such as changing “The image shows…” to “The image and audio show…”, to accurately match the inputs.
@@ 2.3 Evaluation
Our safety evaluation framework first assesses safety performance on individual modality variations, then evaluates the cross-modal safety consistency across all tested variations.
@@ 2.3.1 Safety Evaluation on Individual Modality Variations
Overview. LLM-as-a-judge (Gu et al., 2024) is used to evaluate three aspects: whether the model understands the question, if its output contains harmful content, and if its outputs indicates a refusal to answer. From these, we compute Conditional Attack Success Rate (C-ASR): the probability of harmful content when understood, and Conditional Refusal Rate (C-RR): the probability of refusal when understood. Finally, we derive a Safety-score from C-ASR and C-RR.
Motivation for Comprehension Evaluation. As previously noted, OLLMs can process inputs combining various modalities, with more complex combinations presenting greater comprehension challenges that may lead to irrelevant responses. Our testing revealed that models failing to generate harmful responses due to inability to understand multimodal inputs can artificially inflate safety scores when such interactions are included in evaluation. Examples are provided in Appendix E.2.
C-ASR and C-RR. The metrics are calculated as follows:
$\text{C-ASR}=\frac{N_{\text{safe}=\text{true}\land\text{understand}=\text{true}}}{N_{\text{understand}=\text{true}}}\quad,\quad\text{C-RR}=\frac{N_{\text{refuse}=\text{true}\land\text{understand}=\text{true}}}{N_{\text{understand}=\text{true}}}$ (1)
@@
where $\lambda$ is a weighting parameter, set to 0.5 in the experiments. The Safety-score ranges from 0 to 1, where higher values indicate better safety performance. By design, higher C-ASR lowers the score, while higher C-RR raises it. This prioritizes C-ASR in safety checks, as harmful content is the main concern. Refusing to respond is a secondary sign of stronger safety awareness in the model.